OpenStudy (anonymous):

Write the standard form of the equation of the line parallel to the graph of x – 2y – 6 = 0 and passing through C(0, 1).

OpenStudy (ajprincess):

First covert the given equation into slope intercept form y=mx+c m-slope c-y-intercept x-2y-6=0 x-6=2y (x-6)/2=y x/2-3=y slope is same for parallel lines. y-y_1=m(x-x_1) y-1=m(x-0) Can u do nw @stevenc?

OpenStudy (anonymous):

yes. the answer is y=1/2x+1

OpenStudy (anonymous):

but in standard form I think it is -1/2x+y=1

OpenStudy (anonymous):

@ajprincess

OpenStudy (ajprincess):

ya u r right. bt it wud be better if u make the coefficient of x positive.

OpenStudy (ajprincess):

you can do that by multiplying the whole equation by -1.